OpenStudy (experimentx):

Evaluate (copied): \[ \int\cot^{-1} (x^2+x+1)dx \]

OpenStudy (anonymous):

\[\int\cot^{-1}\left(x^2+x+1\right)dx\\ \int\cot^{-1}\left(\left(x+\frac{1}{2}\right)^2+\frac{3}{4}\right)dx\] \[u=x+\frac{1}{2}\\ du=dx\] \[\int\cot^{-1}\left(u^2+\frac{3}{4}\right)du\] Integrating by parts, let \[\begin{matrix}f=\cot^{-1}\left(u^2+\frac{3}{4}\right)& &dg=du\\ df=-\frac{2u}{\left(u^2+\frac{3}{4}\right)^2+1}du& &g=u\end{matrix}\] \[u\cot^{-1}\left(u^2+\frac{3}{4}\right)+\int\frac{2u^2}{\left(u^2+\frac{3}{4}\right)^2+1}du\] I'm not sure how to approach this next integral, though... I was thinking a sub like \[u=\sqrt t, \text{ then maybe }\\t+\frac{3}{4}=\tan y,\] but it doesn't look like it'll all work out well. Wolframalpha suggests partial fractions, but I don't see how that would work.

OpenStudy (experimentx):

try integration by parts ...

OpenStudy (anonymous):

You mean on the newest integral, right?

OpenStudy (experimentx):

no .. on the original one. and try to use the property of arctan ... this is freaking tricky, otherwise it would get complicated.

OpenStudy (anonymous):

Ah, that makes it much simpler.

OpenStudy (anonymous):

\[\cot ^{-1}x=\tan ^{-1}\frac {1}{x}\]\[\int \tan^{-1}\frac{1}{x}dx\]

OpenStudy (anonymous):

not sure is this right substitution

OpenStudy (anonymous):

Yep, that's how it looks it will start off.

OpenStudy (experimentx):

yep ... that's one!! there is also another property of arctan ... related to addition.

OpenStudy (anonymous):

\[\int \tan^{-1}\frac{1}{x^2+x+1}dx\]

OpenStudy (anonymous):

\[\tan^2 x+1=\cot ^2x\]???

OpenStudy (anonymous):

That's not the right identity.

OpenStudy (anonymous):

tan^2x+1=sec^2x

OpenStudy (experimentx):

No not that one. \[ \arctan a + \arctan b = ?\]

OpenStudy (anonymous):

pi/2??

OpenStudy (anonymous):

a and b related how

OpenStudy (experimentx):

no .. \[ \arctan a + \arctan b = \arctan \left( \frac{a+b}{1-ab}\right)\]

OpenStudy (experimentx):

see if you can apply there.

OpenStudy (anonymous):

from \[\tan(A+B)=\frac{\tan A+\tan B}{1-\tan A*\tan B}\]

OpenStudy (anonymous):

why dont we take f=tan 1/x^2+x+1 g=1\[I=x(\tan^{-1}\frac{1}{x^2+x+1})-\int\limits \frac{1}{1+(x^2+x+1)^2}*x\]

OpenStudy (anonymous):

g'=1 g=x

OpenStudy (experimentx):

yeah, the identity follows from that formula ... but, see if can make it work ... that's terrible ...

OpenStudy (anonymous):

you mean my attempt above is terrible or the identity

OpenStudy (experimentx):

the attempt ... i warned in the my second post.

OpenStudy (experimentx):

how will you break down this ... \[ \int\limits \frac{x}{1+(x^2+x+1)^2} dx\]

OpenStudy (anonymous):

\[ \frac{1}{1+(x^2+x+1)^2}+\frac{x-1}{1+(x^2+x+1)^2}\] can it works

OpenStudy (anonymous):

is this partial fract

OpenStudy (experimentx):

yeah ... how will do you the first one?

OpenStudy (experimentx):

also it's -1 ... 1+x, but the results are same.

OpenStudy (anonymous):

first one???

OpenStudy (experimentx):

 this does not look nice.